OpenStudy (anonymous):

limit question. will give medal. lim (x-1) / [(sqrt of 2x-1) -1] x-->1 The answer is 1. I have the work as well. But, I don't know how they got to the last step. so basically taking the conjugate of the denominator and putting it at the top. I dont understand how they simplified the denominator to cancel out common terms.

OpenStudy (anonymous):

@phi @bibby @PaxPolaris @iambatman @myininaya

OpenStudy (anonymous):

grabbing something to eat, will be back in 20 mins. please help break up the steps after conjugating!

OpenStudy (freckles):

rationalize the bottom

OpenStudy (freckles):

oh so you know to multiply by bottom's conjugate on top and bottom

OpenStudy (freckles):

\[\frac{x-1}{\sqrt{2x-1}-1} \cdot \frac{\sqrt{2x-1}+1}{\sqrt{2x-1}+1} =\frac{(x-1)(\sqrt{2x-1}+1)}{(2x-1)-1}\]

OpenStudy (jdoe0001):

2x -2 when setting x =1, would give 0 though

OpenStudy (anonymous):

hey! @freckles @jdoe0001

OpenStudy (anonymous):

I understand up to that point @freckles ...but after that the denominator becomes (2x-1) -1....which further simplifies to 2(x-1)....I dont get that part

OpenStudy (freckles):

i left you to simplify the bottom there

OpenStudy (freckles):

and then factor the bottom to see a common factor that is on both top and bottom

OpenStudy (anonymous):

I don't get how the denominator gets to 2(x-1)

OpenStudy (anonymous):

yes, the common factor is suppose to be (x-1)...but I dont get it

OpenStudy (freckles):

2x-1-1 2x-2 2(x-1)

OpenStudy (anonymous):

omgd you are right!

OpenStudy (freckles):

are you saying you don't get how (sqrt(2x-1)-1)*(sqrt(2x-1)+1)=2x-1-1?

OpenStudy (freckles):

oh ok

OpenStudy (anonymous):

no i didnt get that simplifying part...thankyou! I get how the answer is one now!

OpenStudy (freckles):

yeah 1 is correct :p